### v1

@@ 3.1 Monodromy in vineyards: a geometric viewpoint
We next drop the assumption that the vines do not touch the diagonal, but still assume that none of the persistence diagrams contain points with higher multiplicity. To make sense of the other definition below and to provide an alternative formulation to the one above, we need the following observation: Let $\tilde{\mathcal{V}}_{\max}(t)$ be a vine defined on its maximal domain $(t_{\min},t_{\max})$, where $t_{\min},t_{\max}\in\mathbb{R}\cup\{\pm\infty\}$, that contains the point $\mathcal{V}(0)$. More formally $\tilde{\mathcal{V}}_{\max}(t)$ is the restriction to the maximal open interval such that the lift of the curve $\gamma:\mathbb{R}\to\mathbb{R}/2\pi\mathbb{Z}:t\mapsto t\;\textrm{mod}\;2\pi$, contains no limit points on the diagonal for this open interval.
@@
If we are given a vine $\mathcal{V}(t)$ defined on $[0,2\pi]$ we call the vine $\tilde{\mathcal{V}}_{\max}(t)$ defined on its maximal domain that coincides with $\mathcal{V}(t)$ on the interval $[0,2\pi]$, the maximal extension of $\mathcal{V}(t)$. If the domain of a maximal extension is finite, then we say that the vine has a finite maximal extension.
@@
If the number of points in each persistence diagram is finite (as we assume), then either both $t_{\min}$ and $t_{\max}$ are (plus/minus) infinite or neither of them are. This is clear because if one of them is infinite we must return to the same point in the persistence diagram after some time $2\pi k$ (since the number of points in the persistence diagram is finite), in which case $\tilde{\mathcal{V}}(t)$ is periodic and is defined on (,)($\infty$,$\infty$)( - ∞ , ∞ ).
@@
Remark 3.2.
Assume that there are no points of higher multiplicity in $\mathscr{D}_{l}(d(\cdot,\gamma(t))_{\mathcal{M}})$, for all $\gamma(t)$ and by extension the vines are non-intersecting curves. Assume moreover that all of the vines are disjoint from the diagonal. The vine $\mathcal{V}(t)$ exhibits monodromy of order $k$ if and only if $k$ is the smallest positive integer such that $\tilde{\mathcal{V}}_{\max}(0)=\tilde{\mathcal{V}}_{\max}(2\pi k)$, where $\tilde{\mathcal{V}}_{\max}(t)$ is the maximal extension of $\mathcal{V}(t)$.
Figure 7: A schematic overview of how we find monodromy in the vineyard. The gray part of the knot does not contribute to monodromy, and is not shown in the vineyard.
In this setting, that is where we allow vines to contain points in the diagonal (as limit points), we make the following definition:
Definition 3.3.
Assume that there are no points of higher multiplicity in $\mathscr{D}_{l}(d(\cdot,\gamma(t))_{\mathcal{M}})$, for all $\gamma(t)$ and by extension the vines are non-intersecting curves. We’ll follow the same notation as above.
Individual vine Let $\mathcal{V}(t)$ be the vine and $\tilde{\mathcal{V}}_{\max}(t)$ is its maximal extension. If its maximal domain is (,)($\infty$,$\infty$)( - ∞ , ∞ ), then the vine does not touch the diagonal and Definition 3.1 and the reformulation in Remark 3.2 both apply. Let us now assume that $t_{\min},t_{\max}$ are finite. By reparametrizing we can assume (without loss of generality) that $t_{\min}=0$. The order of monodromy is now defined as the smallest positive integer $k$, such that $t_{\max}<2\pi k$. As before, we’ll say that the monodromy is trivial if $k=1$.
We’ll also define the completion along the diagonal of a vine $\mathcal{V}$. Let $\tilde{\mathcal{V}}_{\max}:(t_{\min},t_{\max})\to\mathscr{D}_{l}(d(\cdot,\gamma(t))_{\mathcal{M}})$ be the maximal extension of $\mathcal{V}$, and write
@@
for the two limit points of the vine on the diagonal. Let $k$ be the order of monodromy of $\mathcal{V}$ we define the completing diagonal vine $\mathcal{V}_{D}$ as
$\mathcal{V}_{D}:$ $(t_{\max},t_{\min}+2\pi k)\to\mathscr{D}_{l}(d(\cdot,\gamma(t))_{\mathcal{M}})$
@@
The start and end points of the concatenation $\mathcal{V}_{D}\circ\tilde{\mathcal{V}}_{\max}$ coincide and therefore by identifying $t_{\min}$ and $t_{\min}+2\pi k$ we can consider this to be a map on $\mathbb{R}/2\pi k\mathbb{Z}$. The set $\mathbb{R}/2\pi k\mathbb{Z}$ can be viewed as $k$ cover of the circle $\mathbb{R}/2\pi\mathbb{Z}$, where we identify $\mathbb{R}/2\pi\mathbb{Z}$ with the loop $\gamma$. Composing with this cover map gives a map $\mathcal{V}_{C}:\gamma(t)\mapsto\mathscr{D}_{l}(d(\cdot,\gamma(t))_{\mathcal{M}})$, which exhibits monodromy in the way we defined above. We call $\mathcal{V}_{C}$ the completion of the vine $\mathcal{V}$.
Vineyard We now call the vines whose maximal extension do not intersect the diagonal (or equivalently those whose maximal domain is (,)($\infty$,$\infty$)( - ∞ , ∞ )) non-rooted vines, while we call those that intersect the diagonal (or equivalently those whose maximal domain has finite length) rooted vines. We write $k_{1},\dots,k_{n}$ for the orders of monodromy of non-rooted vines and write $l_{1},\dots,l_{m}$ for the orders of monodromy of the rooted vines. Let $l_{\max}=\max\{l_{1},\dots,l_{m}\}$, then we define the order of monodromy of the vineyard as the smallest $k$ that is a common multiple of $k_{1},\dots,k_{n}$ and is larger than $l_{\max}$. We also call $k$ the common order of monodromy of all the vines.
@@
We stress that there may be self-intersections on the diagonal of the completed vineyard. We note again here that if the vineyards are non-generic, the situation is significantly more complex as vines can intersect [28]; we will not consider the non-generic case any further in this work.
3.2 Braids and Morse theory
We say that a closed braid $B$ is $(\epsilon,R)$-embedded in $\mathbb{R}^{3}$ if the closed braid is contained in the $\epsilon$-thickening of the circle $C_{h}(0,R)$ of radius $R$ contained in the horizontal plane. We will tacitly assume that $\epsilon\leq R$.
We say that the maximal angle that a strand of a braid makes with the horizontal direction is the maximal braid angle $\theta_{B}$. For a closed braid we say that the maximal braid angle $\theta_{B}$ is the angle that the tangent line to a strand at a point $p$ in the braid makes with the normal of the plain $P$ defined by $p$ and the $z$-axis.
We can consider a small [^2] The offset should be small enough such that no self-intersection, nor intersections with the $z$-axis occur. $(l+1)$-dimensional $\alpha$-offset $\mathcal{M}$ of a closed braid $B$. This offset is defined as follows: For $l=0$, we do not add an offset. For $l\geq 1$, we consider the embedding $\mathcal{T}\times 0$ in $\mathbb{R}^{3}\times\mathbb{R}^{l-1}$, where $\mathbb{R}^{3}$ is the space that contains the standard embedding of the torus, in the sense of Definition 2.1. We take $\mathcal{M}$ to be the offset of the braid $B\times 0$ in $\mathbb{R}^{3}\times\mathbb{R}^{l-1}$, that is the boundary of $(\mathcal{T}\times 0)\oplus B(0,\alpha)$, where $\oplus$ denotes the Minkowski sum. The resulting manifold $\mathcal{M}$ can then be embedded in $\mathbb{R}^{d}$ where the $\mathbb{R}^{3}$ that contained $\mathcal{T}$ corresponds to the first three coordinates. The maximal braid angle in this context is defined as follows: Let $p\in\mathcal{M}$ and $T_{p}\mathcal{M}$ be its tangent space, then the maximal braid angle is the angle between the normal $n$ of the hyperplane $P\times\mathbb{R}^{d-3}$, where $P$ is the plane defined by the $z$-direction and the point $p$.
Lemma 3.4.
Let $B$ be $(\epsilon,R)$-embedded closed braid and suppose that $\mathcal{M}$ is its $(l+1)$-dimensional $\alpha$-offset, with braid angle $\theta_{B}$. If $p\in\mathbb{R}^{d}$ satisfies $d(p,C_{h}(0,R))\leq\eta$, $C_{h}(0,R)$ is parametrized by $s(\theta)$, and the closest point $p^{\prime}$ of $p$ on $C_{h}(0,R)$ satisfies $p^{\prime}=s(0)$. Then the function $x\mapsto d(x,p)_{\mathcal{M}}$ has no critical points at the closest point $\beta(\theta)$ on $B$ to $x$ to as long as
$\frac{\theta}{2}+\theta_{B}+\arcsin\frac{\epsilon}{2R\sin(\frac{\theta}{2})-\eta}+\arcsin\frac{\eta}{2R\sin(\frac{\theta}{2})}<\frac{\pi}{2}.$ (4)
This implies in particular that there is no topological change as long as (4) is satisfied.
We refer to Figure 10 for a situation sketch.
The proof of this lemma depends on one of the Morse theorems, see e.g. [20, Theorem 3.1]:
Theorem 3.5.
Suppose $f$ is a smooth real-valued function on $\mathcal{M}$, $a<b$, $f^{-1}[a,b]$ is compact, and there are no critical values between $a$ and $b$. Then $M^{a}$ a is diffeomorphic to $M^{b}$, and $M^{b}$ deformation retracts onto $M^{a}$.
Proof 3.6 (of Lemma 3.4).
We write $\beta(\theta)$ for a parametrization of $B$ according to the angle of the circle $C_{h}(0,R)$. We parametrize the circle as $s(\theta)=(R\sin\theta,R\cos\theta,0,\dots,0)$. Using this notation we note that the gradient of $d(y,p)_{\mathcal{M}}$ is zero at $y=x$ if for its closest point on $B$, that is $\beta(\theta)$, we have that $\langle{\beta}^{\prime},\beta-p\rangle=0$, where $\beta^{\prime}$ denotes the derivative of $\beta$ with respect to $\theta$. This is equivalent to
@@ 3.2 Braids and Morse theory
as can be seen from Figure 8.
Figure 8: The angle estimate for $\angle\beta-p,s-p$.
@@
$|\angle({\beta}^{\prime},\beta-p)-\angle({s}^{\prime},s-p)|\leq\theta_{B}+\arcsin\frac{\epsilon}{|s-p|}.$ (5)
Let us now write $d=|s-p|$, $p^{\prime}$ for the closest point projection of $p$ on $C_{h}(0,R)$ and $\omega=\angle({s}^{\prime},s-p^{\prime})$. By reparameterization we can assume that $p^{\prime}=s(0)$. With this assumption we see by the construction in Figure 9 that $\omega=\frac{\theta}{2}$. Moreover we have that $|p^{\prime}-s|=2R\sin(\frac{\theta}{2})$
Figure 9: The construction for the angle $\omega=\angle({s}^{\prime},s-p^{\prime})$.
By the same argument as given in Figure 8 we have that
@@
so that together with (5) we find that
@@
$\leq\theta_{B}+\arcsin\frac{\epsilon}{|s-p^{\prime}|-\eta}+\arcsin\frac{\eta}{|s-p^{\prime}|}$ (by the triangle inequality)
@@
then $\angle({\beta}^{\prime},\beta-p)\neq\pi/2$ and hence there are no critical points. The fact that there is no topological change follows from Theorem 3.5.
Remark 3.7.
The important conclusion from the bound of Lemma 3.4 is that by choosing $\epsilon<\eta\simeq\theta_{B}$ small compared to $\min\{R,\pi\}$, the bound (4) is satisfied as long as $\theta_{B}\ll\theta<\pi-4\theta_{B}$.
Lemma 3.8.
Let $B$ be an oriented $(\epsilon,R)$-embedded, $\delta$-circular closed braid, with braid angle $\theta_{B}$. If $p\in\mathbb{R}^{d}$ satisfies $d(p,C_{h}(0,R))\leq\eta$, $C_{h}(0,R)$ is parametrized by $s(\theta)$, and the closest point $p^{\prime}$ of $p$ on $C_{h}(0,R)$ satisfies $p^{\prime}=s(0)$.
and $\epsilon<\eta\simeq\theta_{B}$ are small compared to $\min\{R,\pi\}$, then $d(\cdot,p)|_{B}$ has $2n$ critical points, $n$ maxima and $n$ minima. Let $l\geq 1$. If $\mathcal{M}$ is a $(l+1)$-dimensional $\alpha$ offset $\mathcal{M}$ of the same type of braid satisfying the same conditions then $d(\cdot,p)|_{\mathcal{M}}$ has $4n$ critical points, $n$ maxima and $n$ minima and $2n$ saddle points of index $l$.
Proof 3.9.
Because the conditions of Lemma 3.4 are satisfied we know that there are no Morse critical points unless $\theta\simeq 0$ or $\theta\simeq\pi$ so we focus on establishing that there is only one critical point per strand at $\theta\simeq 0$ or $\theta\simeq\pi$ for $B$ and or two per strand at $\theta\simeq 0$ or $\theta\simeq\pi$ in the case of $\mathcal{M}$.
The proof for the braid case is the difficult step, and we will see below that the statement for $\mathcal{M}$ follows immediately. The idea of the proof is the following: Because the closed braid is $\delta$-circular its parametrization $\beta$ satisfies $\left|\frac{\beta(t)}{R^{2}}+\ddot{\beta}(t)\right|<\delta$. This means that $\beta$ is forced to turn inward towards the centre of $C_{h}(0,R)$, following the circle $C_{h}(0,R)$, where 00 is the origin of Euclidean space. This in turn implies that $\beta$ cannot satisfy $\frac{d}{dt}|\beta-p|^{2}(t)=0$ for two times $t$ that are relatively close. In other words the curve cannot be tangent to some sphere (not necessarily of the same radius) centred at $p$.
The way that we establish that $\frac{d}{dt}|\beta-p|^{2}(t)$ cannot be zero for two nearby values is by establishing that if $\frac{d}{dt}|\beta-p|^{2}(t)=0$ then
@@
$=2\langle\ddot{\beta},\beta-p\rangle+2$ (because $|\dot{\beta}|=1$)
@@
$=-2\left\langle\frac{\beta}{R^{2}},\beta\right\rangle+2\left\langle\Delta,\beta-p\right\rangle+2+2\left\langle\frac{\beta}{R^{2}},p\right\rangle$ (6)
We can now examine the first three terms in (6):
@@
    $-2\left(1+\frac{\epsilon}{R}\right)^{2}\leq-2\left\langle\frac{\beta}{R^{2}},\beta\right\rangle\leq-2\left(1-\frac{\epsilon}{R}\right)^{2},$
@@
    $-2-\frac{6\epsilon}{R}\leq-2\left\langle\frac{\beta}{R^{2}},\beta\right\rangle\leq-2+\frac{6\epsilon}{R}.$
@@
This implies that these three terms are close to zero, i.e.
@@
For the second part, note that there is a one-to-one correspondence between pairs of critical points of the distance function to a fixed point on an offset and the critical points of the distance function to the same fixed point on the curve $\beta$ itself.
Given these $4n$ critical points, we can now consider the persistence diagram that results. Although we work with extended persistence throughout this paper in order to avoid points at infinity and establish a perfect pairing of critical points, we note that in fact for the purposes of establishing monodromy in our knot offset, it suffices to restrict our attention to the behavior of the ordinary points as well as one single extended point, all above the diagonal, as those that contribute to the construction of knot or link as mentioned in Theorem 1.2 are born and die ‘on the way up’, that is in the first line of the sequence in (2).
Corollary 3.10.
Under the same assumptions as in Lemma 3.8, we have that the maxima and minima of $d(\cdot,p)|_{B}$ and $d(\cdot,p)|_{\mathcal{M}}$ as well as the saddle points of $d(\cdot,p)|_{\mathcal{M}}$ can be divided into two groups, one group occurring at low values and corresponding to $H_{0}$ births of cycles in the persistence diagram and one group occurring at high values and corresponding to $H_{0}$ deaths in the persistence diagram. For $H_{l}$ the situation is identical except for one change, namely that a single birth occurs at a high value. In ‘ordinary’ persistence theory this cycle lives forever, while in extended persistence it dies at the global minimum, and in fact lies below the diagonal.
Proof 3.11.

### v2

@@ B Monodromy touching the diagonal
If we drop the assumption that the vines do not touch the diagonal, but still assume that none of the persistence diagrams contain points with higher multiplicity, it is possible to formulate a notion of monodromy that is more equivalent to that contained in prior work [4]. To make sense of this definition given below and to provide an alternative formulation to the one in our main paper, we need the following observation: Let $\tilde{\mathcal{V}}_{\max}(t)$ be a vine defined on its maximal domain $(t_{\min},t_{\max})$, where $t_{\min},t_{\max}\in\mathbb{R}\cup\{\pm\infty\}$, that contains the point $\mathcal{V}(0)$. More formally $\tilde{\mathcal{V}}_{\max}(t)$ is the restriction to the maximal open interval such that the lift of the curve $\gamma:\mathbb{R}\to\mathbb{R}/2\pi\mathbb{Z}:t\mapsto t\;\textrm{mod}\;2\pi$, contains no limit points on the diagonal for this open interval.
@@
If we are given a vine $\mathcal{V}(t)$ defined on $0,2\pi]$ we call the vine $\tilde{\mathcal{V}}_{\max}(t)$ defined on its maximal domain that coincides with $\mathcal{V}(t)$ on the interval $0,2\pi]$, the maximal extension of $\mathcal{V}(t)$. If the domain of a maximal extension is finite, then we say that the vine has a finite maximal extension.
@@
If the number of points in each persistence diagram is finite (as we assume), then either both $t_{\min}$ and $t_{\max}$ are (plus/minus) infinite or neither of them are. This is clear because if one of them is infinite we must return to the same point in the persistence diagram after some time $2\pi k$ (since the number of points in the persistence diagram is finite), in which case $\tilde{\mathcal{V}}(t)$ is periodic and is defined on $(-\infty,\infty)$.
@@
Remark B.1.
Assume that there are no points of higher multiplicity in $\mathscr{D}_{l}(d_{\mathbb{E}}(\cdot,\gamma(t))|_{\mathcal{M}})$, for all $\gamma(t)$ and by extension the vines are non-intersecting curves. Assume moreover that all of the vines are disjoint from the diagonal. The vine $\mathcal{V}(t)$ exhibits monodromy of order $k$ if and only if $k$ is the smallest positive integer such that $\tilde{\mathcal{V}}_{\max}(0)=\tilde{\mathcal{V}}_{\max}(2\pi k)$, where $\tilde{\mathcal{V}}_{\max}(t)$ is the maximal extension of $\mathcal{V}(t)$.
In this setting, that is, where we allow vines to contain points in the diagonal (as limit points), the goal again is to glue together in a way that we have the exhibited periodicity of monodromy as an integer. More formally:
Definition B.2.
Assume that there are no points of multiplicity higher than $1$ in $\mathscr{D}_{l}(d_{\mathbb{E}}(\cdot,\gamma(t))|_{\mathcal{M}})$, for all $\gamma(t)$ and by extension the vines are non-intersecting curves. We’ll follow the same notation as above.
Individual vine monodromy Let $\mathcal{V}(t)$ be the vine and $\tilde{\mathcal{V}}_{\max}(t)$ its maximal extension. If its maximal domain is $(-\infty,\infty)$, then the vine does not touch the diagonal and Definition 3.2 and the reformulation in Remark B.1 both apply. Let us now assume that $t_{\min},t_{\max}$ are finite. By reparametrizing we can assume (without loss of generality) that $t_{\min}=0$. The order of monodromy is now defined as the smallest positive integer $k$, such that $t_{\max}<2\pi k$. As before, we’ll say that the monodromy is trivial if $k=1$.
We’ll also define the completion along the diagonal of a vine $\mathcal{V}$. Let $\tilde{\mathcal{V}}_{\max}:(t_{\min},t_{\max})\to\mathscr{D}_{l}(d_{\mathbb{E}}(\cdot,\gamma(t))|_{\mathcal{M}})$ be the maximal extension of $\mathcal{V}$, and write
@@
for the two limit points of the vine on the diagonal. Letting $k$ be the order of monodromy of $\mathcal{V}$, we define the completing diagonal vine $\mathcal{V}_{D}$ as
$\mathcal{V}_{D}:$ $(t_{\max},t_{\min}+2\pi k)\to\mathscr{D}_{l}(d_{\mathbb{E}}(\cdot,\gamma(t))|_{\mathcal{M}})$
@@
The start and end points of the concatenation $\mathcal{V}_{D}\circ\tilde{\mathcal{V}}_{\max}$ coincide and therefore by identifying $t_{\min}$ and $t_{\min}+2\pi k$ we can consider this to be a map on $\mathbb{R}/2\pi k\mathbb{Z}$. The set $\mathbb{R}/2\pi k\mathbb{Z}$ can be viewed as a $k$ cover of the circle $\mathbb{R}/2\pi\mathbb{Z}$, where we identify $\mathbb{R}/2\pi\mathbb{Z}$ with the loop $\gamma$. Composing with this cover map gives a map $\mathcal{V}_{C}:\gamma(t)\mapsto\mathscr{D}_{l}(d_{\mathbb{E}}(\cdot,\gamma(t))|_{\mathcal{M}})$, which exhibits monodromy in the way we defined above. We call $\mathcal{V}_{C}$ the completion of the vine $\mathcal{V}$.
Vineyard monodromy We now call the vines whose maximal extension do not intersect the diagonal (or equivalently those whose maximal domain is $(-\infty,\infty)$) non-rooted vines, while we call those that intersect the diagonal (or equivalently those whose maximal domain has finite length) rooted vines. We write $k_{1},\dots,k_{n}$ for the orders of monodromy of non-rooted vines and write $l_{1},\dots,l_{m}$ for the orders of monodromy of the rooted vines. Let $l_{\max}=\max\{l_{1},\dots,l_{m}\}$, then we define the order of monodromy of the vineyard as the smallest $k$ that is a common multiple of $k_{1},\dots,k_{n}$ and is larger than $l_{\max}$. We also call $k$ the common order of monodromy of all the vines.
@@
We stress that there may be self-intersections on the diagonal of the completed vineyard. We note again here that if the vineyards are non-generic, the situation is significantly more complex, as vines can intersect [47]; we will not consider the non-generic case any further in this work, but it may be of interest for future study.
C Omitted Proofs
We now present more details on the technical proofs that were omitted in Section 4.1.
C.1 Proof of Lemma 4.1
The proof of this lemma depends on one of the Morse theorems, see e.g. [32, Theorem 3.1]:
Theorem C.1 (Morse).
Suppose $f$ is a smooth real-valued function on $\mathcal{M}$, $a<b$, $f^{-1}[a,b]$ is compact, and there are no critical values between $a$ and $b$. Then $\mathcal{M}^{a}$ is diffeomorphic to $\mathcal{M}^{b}$, and $\mathcal{M}^{b}$ deformation retracts onto $\mathcal{M}^{a}$.
Proof 3 (of Lemma 4.1).
We write $\beta(\theta)$ for a parametrization of $B$ according to the angle of the circle $C_{h}(0,R)$. We parametrize the circle as $s(\theta)=(R\sin\theta,R\cos\theta,0,\dots,0)$. Using this notation, we note that the gradient of $d(y,p)_{\mathcal{M}}$ is zero at $y=x$ if for its closest point on $B$, that is, $\beta(\theta)$, we have that $\langle{\beta}^{\prime},\beta-p\rangle=0$, where $\beta^{\prime}$ denotes the derivative of $\beta$ with respect to $\theta$. This is equivalent to
@@ C Omitted Proofs
as can be seen from Figure C.1.
Figure C.1: The angle estimate for $\angle\beta-p,s-p$.
@@
(C.1) $|\angle({\beta}^{\prime},\beta-p)-\angle({s}^{\prime},s-p)|\leq\theta_{B}+\arcsin\frac{\epsilon}{|s-p|}.$
Let us now write d=|sp|d=|s-p|, $p^{\prime}$ for the closest point projection of $p$ on $C_{h}(0,R)$ and $\omega=\angle({s}^{\prime},s-p^{\prime})$. By reparameterization we can assume that $p^{\prime}=s(0)$. With this assumption we see by the construction in Figure C.2 that $\omega=\angle({s}^{\prime},s-p^{\prime})=\frac{\theta}{2}$. Moreover we have that $|p^{\prime}-s|=2R\sin(\frac{\theta}{2})$
Figure C.2: The construction for the angle $\omega=\angle({s}^{\prime},s-p^{\prime})$.
By the same argument as that given in Figure C.1 we have
@@
so that together with (C.1) we find that
@@
(by the triangle inequality) $\leq\theta_{B}+\arcsin\frac{\epsilon}{|s-p^{\prime}|-\eta}+\arcsin\frac{\eta}{|s-p^{\prime}|}$
@@
then $\angle({\beta}^{\prime},\beta-p)\neq\pi/2$ and hence there are no critical points. The fact that there is no topological change follows from Theorem C.1.
C.2 Proof of Lemma 4.3
Proof 4.
Because the conditions of Lemma 4.1 are satisfied, we know that there are no Morse critical points unless $\theta\simeq 0$ or $\theta\simeq\pi$ so we focus on establishing that there is only one critical point per strand at $\theta\simeq 0$ or $\theta\simeq\pi$ for $B$ and or two per strand at $\theta\simeq 0$ or $\theta\simeq\pi$ in the case of $\mathcal{M}$.
The proof for the braid case is the difficult step, and we will see below that the statement for $\mathcal{M}$ follows immediately. The idea of the proof is the following: Because the closed braid is $\delta$-circular its parametrization $\beta$ satisfies $\left|\frac{\beta(t)}{R^{2}}+\ddot{\beta}(t)\right|<\delta$. This means that $\beta$ is forced to turn inward towards the centre of $C_{h}(0,R)$, following the circle $C_{h}(0,R)$, where 0 is the origin of Euclidean space. This in turn implies that $\beta$ cannot satisfy $\frac{d}{dt}|\beta-p|^{2}(t)=0$ for two times $t$ that are relatively close. In other words, the curve cannot be tangent to some sphere (not necessarily of the same radius) centred at $p$.
The way we establish that $\frac{d}{dt}|\beta-p|^{2}(t)$ cannot be zero for two nearby values is by establishing that if $\frac{d}{dt}|\beta-p|^{2}(t)=0$ then
@@
(because $|\dot{\beta}|=1$) $=2\langle\ddot{\beta},\beta-p\rangle+2$
@@
(C.2) $=-2\left\langle\frac{\beta}{R^{2}},\beta\right\rangle+2\left\langle\Delta,\beta-p\right\rangle+2+2\left\langle\frac{\beta}{R^{2}},p\right\rangle$
We can now examine the first three terms in (C.2):
@@
    $2\left(1+\frac{\epsilon}{R}\right)^{2}\leq-2\left\langle\frac{\beta}{R^{2}},\beta\right\rangle\leq-2\left(1-\frac{\epsilon}{R}\right)^{2},$
@@
    $2-\frac{6\epsilon}{R}\leq-2\left\langle\frac{\beta}{R^{2}},\beta\right\rangle\leq-2+\frac{6\epsilon}{R}.$
@@
This implies that these three terms are close to zero, and hence
@@
For the second part, note that there is a one-to-one correspondence between pairs of critical points of the distance function to a fixed point on an offset and the critical points of the distance function to the same fixed point on the curve $\beta$ itself. Here minima of $\beta$ correspond to a pair of a minimum and a critical point of index $l$ on $\mathcal{M}$, while maxima correspond to a pair of a maximum and a critical point of index $1$.
C.3 Proof of Corollary 4.4
Proof 5.